OpenStudy (cademcbain):

A landscaping Company charges $100 plus $15 per hour. Another Landscaping Company charges $75 plus $17 per hour. How long is a job that costs the same no matter which company is used

OpenStudy (retireed):

Find where those two lines intersect.

OpenStudy (cademcbain):

then what

OpenStudy (retireed):

The intersection of the two lines is the value of x (hours) where the two jobs cost the same.

OpenStudy (cademcbain):

ok so whats the answer

OpenStudy (cademcbain):

I got to go

OpenStudy (cademcbain):

bye

OpenStudy (retireed):

You tell me ... 

OpenStudy (retireed):

Set the two equations equal to each... 100+15x = 75+17x Solve for x 25 = 2x